### v1

# Is Algorithmic Stability Testable? A Unified Framework under Computational Constraints

Abstract
Algorithmic stability is a central notion in learning theory that quantifies the sensitivity of an algorithm to small changes in the training data. If a learning algorithm satisfies certain stability properties, this leads to many important downstream implications, such as generalization, robustness, and reliable predictive inference. Verifying that stability holds for a particular algorithm is therefore an important and practical question. However, recent results establish that testing the stability of a black-box algorithm is impossible, given limited data from an unknown distribution, in settings where the data lies in an uncountably infinite space (such as real-valued data). In this work, we extend this question to examine a far broader range of settings, where the data may lie in any space—for example, categorical data. We develop a unified framework for quantifying the hardness of testing algorithmic stability, which establishes that across all settings, if the available data is limited then exhaustive search is essentially the only universally valid mechanism for certifying algorithmic stability. Since in practice, any test of stability would naturally be subject to computational constraints, exhaustive search is impossible and so this implies fundamental limits on our ability to test the stability property for a black-box algorithm.
1 Introduction
For data lying in space $\mathcal{X}\times\mathcal{Y}$, an algorithm is a map that maps data sets to fitted regression functions:
$\mathcal{A}:\ \cup_{n\geq 0}(\mathcal{X}\times\mathcal{Y})^{n}\rightarrow\big{\{}\textnormal{measurable functions $\mathcal{X}\rightarrow\mathbb{R}$}\big{\}}.$ (1)
That is, $\mathcal{A}$ takes as input $\mathcal{D}=\{(X_{i},Y_{i})\}_{i\in[n]}$, a data set of any size $n$, and returns a fitted model ${\widehat{f}}_{n}:\mathcal{X}\rightarrow\mathbb{R}$, which maps a feature value $x$ to a fitted value, or prediction, ${\widehat{y}}={\widehat{f}}_{n}(x)$. Algorithmic stability is a notion that quantifies the sensitivity of an algorithm to small changes in the training data $\mathcal{D}$. Since its introduction (Devroye and Wagner,, 1979; Rogers and Wagner,, 1978; Kearns and Ron,, 1997), algorithmic stability has been a central notion in machine learning and statistics. Intuitively, it is a desirable property for learning algorithms—but it also leads to many downstream implications, as well. For instance, it is known that stable algorithms generalize (Kearns and Ron,, 1997; Bousquet and Elisseeff,, 2002; Poggio et al.,, 2004; Shalev-Shwartz et al.,, 2010). It also plays a critical role in a number of different problems, such as model selection (Meinshausen and Bühlmann,, 2010; Shah and Samworth,, 2013; Ren et al.,, 2023), reproducibility (Yu,, 2013), predictive inference (Steinberger and Leeb,, 2023; Barber et al.,, 2021; Liang and Barber,, 2023), selective inference (Zrnic and Jordan,, 2023) and inference for cross-validation error (Austern and Zhou,, 2020; Bayle et al.,, 2020; Kissel and Lei,, 2022).
Thus, verifying the stability of a given algorithm is an important and practical question. Some algorithms can easily be shown to be stable, such as $k$-nearest neighbors (kNN) (Devroye and Wagner,, 1979) and ridge regression (Bousquet and Elisseeff,, 2002). More recently, substantial progress has been made towards understanding the stability for more sophisticated algorithms, see Hardt et al., (2016); Lei and Ying, (2020); Deng et al., (2021) and references therein. However, many complex modern machine learning algorithms exhibit instability empirically, or even if this is not the case, they may be too complex for a theoretical stability guarantee to be possible to obtain. It is common to attempt to stabilize a black-box algorithm by using generic methods such as bootstrap, bagging, or subsampling (Efron and Tibshirani,, 1994; Breiman,, 1996; Elisseeff et al.,, 2005; Chen et al.,, 2022; Soloff et al.,, 2024) so that stability is ensured—but these methods also have their limitations, such as loss of statistical efficiency due to subsampling. Consequently, this motivates the question of whether we can estimate the stability of an algorithm through empirical evaluation, rather than theoretical calculations. This idea is formalized in the work Kim and Barber, (2023) where the authors consider inferring the stability of an algorithm through “black-box” testing without peering inside the construction of the algorithm. Under the setting when $\mathcal{X}$ and $\mathcal{Y}$ spaces have uncountably infinitely many elements and the black-box test can call $\mathcal{A}$ countably infinitely many times, Kim and Barber, (2023) established a fundamental bound on our ability to test for algorithmic stability in the absence of any assumptions on the data distribution and the algorithm $\mathcal{A}$.
However, in many common settings, such as categorical features or a binary response, we may have finite data spaces $\mathcal{X}$ and/or $\mathcal{Y}$. The results of Kim and Barber, (2023), which assume $\mathcal{X}$ and/or $\mathcal{Y}$ to be infinite, cannot be applied to characterize the hardness of testing stability in these settings—and indeed, we would expect the testing problem to be fundamentally easier if there are only finitely many possible values of the data. On the other hand, in practice, we only have a finite amount of computational resources in order to run our test of stability. This leads to a key motivating question: is it possible to construct a black-box test of algorithmic stability in finite data spaces $\mathcal{X}$ and $\mathcal{Y}$, and what role do computational constraints play in this question?
1.1 Overview of contributions
In this paper, we extend the results in Kim and Barber, (2023) to a far broader range of settings and establish fundamental bounds for testing algorithmic stability with computational constraints. Specifically, we provide an explicit upper bound on the power of any universally valid black-box test for testing algorithmic stability in terms of the number of available data, stability parameters, computational budget, and the sizes of the data spaces $\mathcal{X}$ and $\mathcal{Y}$.
Organization.
In Section 2, we will present some background on algorithmic stability and black-box testing framework under computational constraints. Section 3 presents our main results on the limit of black-box tests for algorithmic stability. A brief discussion is provided in Section 4. We prove the main results in Section 5. Some technical details and extensions are deferred to the Appendix.
2 Background and problem setup
In this section, we formalize some definitions and give additional background for studying algorithmic stability via black-box tests. In (1), we have defined the algorithm as a deterministic map given the training data. In many settings, commonly used algorithms might also include some form of randomization—for instance, stochastic gradient descent type methods. To accommodate this, we will expand the definition of an algorithm to take an additional argument,
$\mathcal{A}:\ \cup_{n\geq 0}(\mathcal{X}\times\mathcal{Y})^{n}\times[0,1]\rightarrow\big{\{}\textnormal{measurable functions $\mathcal{X}\rightarrow\mathbb{R}$}\big{\}}.$ (2)
The fitted model ${\widehat{f}}_{n}=\mathcal{A}\big{(}\{(X_{i},Y_{i})\}_{i\in[n]};\xi\big{)}$ is now obtained by running algorithm $\mathcal{A}$ on data set $\{(X_{i},Y_{i})\}_{i\in[n]}$, with randomization provided by the argument $\xi\in[0,1]$, which acts as a random seed. [^1] Throughout this paper, we implicitly assume measurability for all functions we define, as appropriate. For the algorithm $\mathcal{A}$, we assume measurability of the map $\left((x_{1},y_{1}),\ldots,(x_{n},y_{n}),\xi,x\right)\mapsto[\mathcal{A}(\{(x_{i},y_{i})\}_{i\in[n]};\xi)](x)$. Of course, the general definition (2) of a randomized algorithm also accommodates the deterministic (i.e., non-random) setting—we can simply choose an algorithm $\mathcal{A}$ whose output depends only on the data set $\{(X_{i},Y_{i})\}_{i\in[n]}$, and ignores the second argument $\xi$. In this paper, we focus on the setting where $\mathcal{A}$ is symmetric in the training data—that is, permuting the order of the $n$ data points does not change the output of $\mathcal{A}$ (or the distribution of the output, if randomized—see Kim and Barber, (2023) for a formal definition).
2.1 Algorithmic stability and hypothesis testing formulation
There are many different notions of algorithmic stability in the literature (see, e.g., Bousquet and Elisseeff, (2002), Appendix A of Shalev-Shwartz et al., (2010) and Section 2.1.1 of Kim and Barber, (2023) for some comparisons). In this work, we follow Kim and Barber, (2023) and consider testing the following notion of stability:
Definition 1 (Algorithmic stability).
Let $\mathcal{A}$ be a symmetric algorithm. Let $\epsilon\geq 0$ and $\delta\in[0,1)$. We say that $\mathcal{A}$ is $(\epsilon,\delta)$-stable with respect to training data sets of size $n$ from a data distribution $P$—or, for short, the triple $(\mathcal{A},P,n)$ is $(\epsilon,\delta)$-stable—if
$\mathbb{P}\{|{\widehat{f}}_{n}(X_{n+1})-{\widehat{f}}_{n-1}(X_{n+1})|>\epsilon\}\leq\delta,$
where
${\widehat{f}}_{n}=\mathcal{A}\left(\{(X_{i},Y_{i})\}_{i\in[n]};\xi\right),\quad{\widehat{f}}_{n-1}=\mathcal{A}\left(\{(X_{i},Y_{i})\}_{i\in[n-1]};\xi\right),$
and where the data is distributed as $(X_{i},Y_{i})\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$, and the random seed $\xi\sim\textnormal{Unif}[0,1]$ is drawn independently of the data.
In the setting of predictive inference for regression problems, this notion of stability has appeared in Barber et al., (2021) where it was shown that this condition is sufficient for predictive coverage guarantees of the jackknife and jackknife+ methods. Earlier work by Steinberger and Leeb, (2023) also established coverage properties for jackknife under a similar stability property.
Defining a hypothesis test.
Given a fixed $\epsilon\geq 0$ and $\delta\in[0,1)$, we are interested in testing
[…]
with some desired bound $\alpha\in(0,1)$ on the type-I error—that is, a bound on the probability that we falsely certify an algorithm as stable. In this work, we consider the problem of running this hypothesis test with limited access to data from $P$—specifically, we assume that the available data consists of a labeled data set, $\mathcal{D}_{\ell}=\{(X_{i},Y_{i})_{i\in[N_{\ell}]}\}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$, and an unlabeled data set, $\mathcal{D}_{u}=\{X_{N_{\ell}+1},\cdots,X_{N_{\ell}+N_{u}}\}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P_{X}$, where $P_{X}$ is the marginal distribution of $X$ under $P$ (and $\mathcal{D}_{\ell}$ and $\mathcal{D}_{u}$ are sampled independently). With the aim of avoiding placing any assumptions on the algorithm $\mathcal{A}$ and on the distribution $P$, we would like to seek a test ${\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})\in\{0,1\}$ that satisfies the following notion of assumption-free validity:
[…]
2.2 Black-box tests with computational constraints
To understand the difficulty of testing algorithmic stability, we need to formalize the limitations on what information is available to us, as the analyst, for performing the test. For example, if we know the distribution $P$ of the data, then for many algorithms, we might be able to study its stability property analytically. Clearly, this is not the setting of interest, since in practice we cannot rely on assumptions about the data distribution—but in fact, even if $P$ is unknown, analyzing the stability for simple algorithms such as kNN or ridge regression is still tractable (Devroye and Wagner,, 1979; Bousquet and Elisseeff,, 2002). In modern settings, however, a state-of-the-art algorithm $\mathcal{A}$ is typically far more complex and may be too complicated to analyze theoretically. For this reason, we consider the black-box testing framework introduced in Kim and Barber, (2023) which analyzes the stability of algorithm $\mathcal{A}$ via empirical evaluations, rather than theoretical calculations. The intuition is this: if $\mathcal{A}$ is implemented as a function in some software package, our procedure is allowed to call the function, but is not able to examine the implementation of $\mathcal{A}$ (i.e., we cannot read the code that defines the function).
Compared to the definition of black-box test given in Kim and Barber, (2023), here we focus on black-box tests with computational constraints. A high-level illustration of a black-box test with computational constraints is provided in Figure 1; the formal definition is quite technical and we defer it to Appendix A. In particular, the computational budget for training is characterized by $B_{\textnormal{train}}$—the total number of data points that we are allowed to input to the algorithm $\mathcal{A}$, over all calls to $\mathcal{A}$ (for example, when running the test, we can choose to train $\mathcal{A}$ on a data set of size $n$, up to $B_{\textnormal{train}}/n$ many times). Some examples of black-box tests, and counterexamples (i.e., tests that do not follow the black-box framework), can be found in Kim and Barber, (2023, Section 2.3).
Input: labeled data $\mathcal{D}_{\ell}$, unlabeled data $\mathcal{D}_{u}$, algorithm $\mathcal{A}$, seeds $\zeta,\zeta_{1},\zeta_{2},\dots$ Iterate for rounds $r=1,2,\ldots,{\widehat{r}}$, subject to computational constraint $\sum_{r=1}^{{\widehat{r}}}|\mathcal{D}^{(r)}_{\ell}|\leq B_{\textnormal{train}}$ Generate labeled data set $\mathcal{D}^{(r)}_{\ell}$ and random seed $\xi^{(r)}$ Call the algorithm to train a model: ${\widehat{f}}^{(r)}=\mathcal{A}(\mathcal{D}^{(r)}_{\ell};\xi^{(r)})$ Output: ${\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})\in\{0,1\}$
Figure 1: Black-box test with computational constraints. Here in round $r$, the generating process of the new data set $\mathcal{D}_{\ell}^{(r)}$ and random seed $\xi^{(r)}$ for training can depend on the input data and on all the past information from previous rounds—e.g., we may create a new data set by resampling from past data. The input random seeds $\zeta,\zeta_{1},\zeta_{2},\dots\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}\textnormal{Unif}[0,1]$ may be used to provide randomization if desired. The procedure stops at a data-dependent stopping time ${\widehat{r}}$, with the constraint that the total number of training points used in calls to $\mathcal{A}$ cannot exceed the budget $B_{\textnormal{train}}$.
Transparent models or black-box models?
Careful readers may notice that the version of black-box test we presented in Figure 1, is called a black-box test with transparent models by Kim and Barber, (2023). This means that while the training algorithm $\mathcal{A}$ can only be studied empirically via calls to $\mathcal{A}$ (that is, we cannot examine its theoretical properties, or, say, read the code that defines the implementation of $\mathcal{A}$), in contrast the fitted model ${\widehat{f}}^{(r)}$ returned by $\mathcal{A}$ at the $r$th round is transparent to us—we know its analytic expression. This is a less restrictive notion of black-box testing, giving more information to the analyst. In contrast, an even stronger constraint is the notion of a black-box test with black-box models in Kim and Barber, (2023), where the fitted models cannot be studied analytically and instead can only be studied via evaluating the models on finitely many test points.
In the main part of this paper, we will focus on the black-box test with transparent models, since this offers more flexibility to the analyst—and thus, establishing hardness of testing algorithmic stability in this setting is a stronger result. We extend our results to the black-box model setting in Appendix B.
3 Main results
In this section, we present the main results on a unified power upper bound of testing algorithmic stability for computationally constrained black-box tests. First, given any $(\mathcal{A},P,n)$, let us define
$\delta^{*}_{\epsilon}=\mathbb{P}\{|{\widehat{f}}_{n}(X_{n+1})-{\widehat{f}}_{n-1}(X_{n+1})|>\epsilon\},$ (4)
i.e., $\delta^{*}_{\epsilon}$ is the smallest possible value of $\delta^{\prime}$ so that $(\mathcal{A},P,n)$ is $(\epsilon,\delta^{\prime})$-stable.
Theorem 1.
Fix any $\epsilon\geq 0$, $\delta\in[0,1)$. Let ${\widehat{T}}$ be any black-box test with computational budget $B_{\textnormal{train}}$ (as illustrated in Figure 1—see Appendix A for the formal definition). Suppose ${\widehat{T}}$ satisfies the assumption-free validity condition (3) at level $\alpha$.
Then for any $(\mathcal{A},P,n)$ that is $(\epsilon,\delta)$-stable (i.e., $\delta^{*}_{\epsilon}\leq\delta$), the power of ${\widehat{T}}$ is bounded as
$\mathbb{P}\{{\widehat{T}}=1\}\leq\alpha\cdot\min\Bigg{\{}\underbrace{\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{\frac{B_{\textnormal{train}}}{n}}}_{\begin{subarray}{c}\textnormal{computational constraint}\\ \textnormal{due to budget $B_{\textnormal{train}}$}\end{subarray}},\underbrace{\frac{\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{\frac{N_{\ell}}{n}}}{1-\frac{B_{\textnormal{train}}}{|\mathcal{Y}|}\wedge 1}}_{\begin{subarray}{c}\textnormal{statistical constraint due to}\\ \textnormal{limited samples in $\mathcal{Y}$}\end{subarray}},\underbrace{\frac{\left(\frac{1-\delta^{*}_{\epsilon}}{1-\widetilde{\delta}}\right)^{\frac{N_{\ell}+N_{u}}{n}}}{1-\frac{B_{\textnormal{train}}}{|\mathcal{X}|}\wedge 1}}_{\begin{subarray}{c}\textnormal{statistical constraint due to}\\ \textnormal{limited samples in $\mathcal{X}$}\end{subarray}}\Bigg{\}},$ (5)
where $\widetilde{\delta}=\delta(1+\frac{1}{en})\wedge 1$.
In Theorem 1, we can see the power of any universally valid test is bounded by the minimum of three terms. Here we take a closer look at each of the three terms to better understand how this limit on power arises.
  • First term: computational constraint due to budget $B_{\textnormal{train}}$. The first term is low—that is, not much larger than $1$, meaning that the power cannot be much better than random—whenever the ratio $B_{\textnormal{train}}/n$ is small. This is intuitive: this ratio describes the number of times we are able to call $\mathcal{A}$ on training samples of size $n$, and we need this number to be large in order to distinguish between stability (i.e., large perturbations occur with small probability $\delta^{*}_{\epsilon}\leq\delta$) versus instability (i.e., large perturbations occur with probability $>\delta$).
  • Second term: statistical constraint due to limited samples in $\mathcal{Y}$. The second term shows that power is low whenever we have both of the following: the ratio $N_{\ell}/n$ is not too large, and the ratio $B_{\textnormal{train}}/|\mathcal{Y}|$ is close to 0. For the first quantity, a bound on $N_{\ell}/n$ means that we do not have access to ample labeled data—the available data sampled from $P$ provides only $N_{\ell}/n$ many independent data sets of size $n$, so it is not easy to empirically estimate how often instability may occur. For the second ratio, if $B_{\textnormal{train}}/|\mathcal{Y}|\approx 0$ then this means that we do not have sufficient computational budget to allow an exhaustive search over $\mathcal{Y}$ (i.e., we cannot afford to test the behavior of $\mathcal{A}$ on all possible values of $Y$) to compensate for the limited amount of available labeled data.
  • Third term: statistical constraint due to limited samples in $\mathcal{X}$. The interpretation for this term is analogous to the previous term, with the difference that we are given $N_{\ell}+N_{u}$ many draws of $X$ due to the additional unlabeled data set $\mathcal{D}_{u}$.
In particular, the power can be only slightly higher than $\alpha$ (i.e., only slightly better than random), unless all of the three terms in the minimum are large: to have high power, we would need $B_{\textnormal{train}}\gg n$ to enable sufficiently many calls to $\mathcal{A}$, and moreover, to search over $\mathcal{Y}$ space we need either $N_{\ell}\gg n$ or $|\mathcal{Y}|$ to be relatively small, and similarly to search over $\mathcal{X}$ space we need either $N_{\ell}+N_{u}\gg n$ or $|\mathcal{X}|$ to be relatively small. Note the second and third terms are statistical limits, due to limited data (i.e., $N_{\ell}$ and $N_{u}$), but these statistical limits can be overcomed if $|\mathcal{X}|$ and $|\mathcal{Y}|$ are small enough to permit exhaustive search (under the computation constraint $B_{\textnormal{train}}$).
An unlimited budget?
While our result is stated for a black-box test run under a prespecified computational constraint, $B_{\textnormal{train}}$, it is also natural to ask what power might be obtained if we do not specify a budget. In fact, our result applies to this setting as well. Suppose that we require our black-box test ${\widehat{T}}$ to terminate at a finite time (i.e., ${\widehat{r}}$ must be finite), but we do not specify a budget—that is, ${\widehat{r}}$ can be arbitrarily large, and $\sum_{r=1}^{{\widehat{r}}}|\mathcal{D}^{(r)}_{\ell}|$ can be arbitrarily large. In this setting, we can simply apply the results of Theorem 1 while taking $B_{\textnormal{train}}\to\infty$. In particular,
$\lim_{B_{\textnormal{train}}\to\infty}\left(1-\frac{B_{\textnormal{train}}}{|\mathcal{Y}|}\wedge 1\right)=\begin{cases}1,&|\mathcal{Y}|=\infty,\\ 0,&|\mathcal{Y}|<\infty,\end{cases}$
and the same is true for $|\mathcal{X}|$. Therefore, even with an unlimited budget, if we require ${\widehat{T}}$ to return an answer in finite time then Theorem 1 implies the following bounds on power:
$\mathbb{P}\{{\widehat{T}}=1\}\leq\alpha\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{\frac{N_{\ell}}{n}},\textnormal{ if $|\mathcal{Y}|=\infty$; \quad}\mathbb{P}\{{\widehat{T}}=1\}\leq\alpha\left(\frac{1-\delta^{*}_{\epsilon}}{1-\widetilde{\delta}}\right)^{\frac{N_{\ell}+N_{u}}{n}},\textnormal{ if $|\mathcal{X}|=\infty$}.$
Comparison with existing work.
Our result is a strict generalization of Kim and Barber, (2023, Theorem 3) , which proves that power is bounded as $\mathbb{P}\{{\widehat{T}}=1\}\leq\alpha\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{\frac{N_{\ell}}{n}}$ whenever $|\mathcal{Y}|=\infty$. [^2] More precisely, their work assumes that the cardinality of the space $\mathcal{Y}$ is uncountably infinite, but this is because they allow the test ${\widehat{T}}$ to run for countably infinitely many rounds. Since here we require ${\widehat{T}}$ to terminate after finitely many rounds, under our definition of a black-box test, their result would hold also for a countably infinite $\mathcal{Y}$. Of course, this is a special case of our main result, obtained by taking $|\mathcal{Y}|=\infty$ and $B_{\textnormal{train}}\to\infty$. The existing result in their work does not provide a bound on power when $|\mathcal{Y}|<\infty$ (for example, their result does not apply in the common binary classification setting $\mathcal{X}=\mathbb{R}^{d}$ and $\mathcal{Y}=\{0,1\}$), and does not account for the role of the computational budget. The proof of our main result leverages some of the constructions developed in Kim and Barber, (2023)’s proof, but because of the additional considerations on the size of data space as well as the computational constraint, the proof of Theorem 1 requires novel technical tools, as we will see in Section 5.
4 Discussion
In this paper, we provide a unified power upper bound for any universally valid black-box test for testing algorithmic stability under computational constraints. We reveal the computational budget can be the dominant constraint when the budget $B_{\textnormal{train}}$ is small. At the same time, we illustrate that when the data space is finite, if sample size is limited then certifying stability is possible only if the data spaces $\mathcal{X}$ and $\mathcal{Y}$ are small enough to essentially permit an exhaustive search. In this work, we focus on testing $(\epsilon,\delta)$-stability defined in Definition 1; since a wide range of different stability definitions exist in the literature (see, e.g., Bousquet and Elisseeff, (2002)), an interesting open question is whether there is a unified way to treat different notions of stability under the black-box testing framework.
5 Proof of Theorem 1
The theorem can be separated into three claims and we will prove each of these claims separately. Respectively, these three results reflect the bounds on power that are due to a limited number of calls to $\mathcal{A}$ (a computational limit), and to limited $Y$ data and $X$ data (statistical limits).
5.1 Computational limit: bounding power due to limited calls to $\mathcal{A}$
Recall that when we run the test ${\widehat{T}}$, the calls $r=1,\dots,{\widehat{r}}$ to the algorithm $\mathcal{A}$ are carried out with data set $\mathcal{D}^{(r)}_{\ell}$ and with random seed $\xi^{(r)}$ (which may be generated in any fashion). Fixing any subset $R\subset[0,1]$, define $\mathcal{E}_{R}$ as the event that $\xi^{(r)}\not\in R$ for all $r$ such that $|\mathcal{D}^{(r)}_{\ell}|=n$—that is, for all calls $r$ to the algorithm during which the training set size is equal to $n$. Let $\textnormal{Leb}(R)$ denote the Lebesgue measure of $R$. Define also the function $f(\xi)=\mathbb{P}\left(|\mathcal{A}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon\mid\xi\right)$, the conditional probability of an unstable prediction, as a function of the random seed $\xi$ (here the probability is taken with respect to training and test data sampled i.i.d. from $P$). By definition, we have $\mathbb{E}[f(\xi)]=\delta^{*}_{\epsilon}$, where the expected value is taken with respect to $\xi\sim\textnormal{Unif}[0,1]$.
We will need the following lemma:
Lemma 1.
[…]
With this lemma in place, we are ready to bound the power. Fixing any integer $M\geq 1$, standard measure-theoretic arguments (e.g., Lemma 2.5 of Bennett and Sharpley, (1988)) allow us to construct a partition $[0,1]=S_{1}\cup\dots\cup S_{M}$, where $S_{i}$ is measurable and has Lebesgue measure $1/M$, and $\mathbb{E}[f(\xi)\mid\xi\in S_{i}]=\mathbb{E}[f(\xi)]=\delta^{*}_{\epsilon}$, for each $i\in[M]$.
Next, fix any integer $M_{0}\leq M$ with $M_{0}>M\cdot\frac{\delta-\delta_{\epsilon}^{*}}{1-\delta_{\epsilon}^{*}}$. For any subset $I\subset[M]$ with $|I|=M_{0}$, define a region $R_{I}=\cup_{i\in I}S_{i}$, and observe that $\textnormal{Leb}(R_{I})=M_{0}/M>\frac{\delta-\delta_{\epsilon}^{*}}{1-\delta_{\epsilon}^{*}}$ by construction, and $\mathbb{E}[f(\xi)\mid\xi\not\in R_{I}]=\delta^{*}_{\epsilon}$ by construction of the partition $S_{1}\cup\dots\cup S_{M}$. Then by Lemma 1, $\mathbb{P}\{{\widehat{T}}=1;\mathcal{E}_{R_{I}}\}\leq\alpha$. Now we sum over all subsets $I$:
$\sum_{I\subset[M],|I|=M_{0}}\mathbb{P}\{{\widehat{T}}=1;\mathcal{E}_{R_{I}}\}=\mathbb{E}\left[{\mathds{1}}\{{\widehat{T}}=1\}\cdot\!\!\!\!\sum_{I\subset[M],|I|=M_{0}}\!\!\!\!{\mathds{1}}\{\mathcal{E}_{R_{I}}\}\right]\geq\mathbb{P}\{{\widehat{T}}=1\}\cdot{M-\lfloor\frac{B_{\textnormal{train}}}{n}\rfloor\choose M_{0}},$
where the inequality holds as follows: defining $\widehat{I}\subset[M]$ as the set of all indices $i\in[M]$ such that $\{|\mathcal{D}^{(r)}_{\ell}|=n$ and $\xi^{(r)}\in S_{i}\}$ holds for some $r\in\{1,\dots,{\widehat{r}}\}$, we have $\sum_{I\subset[M],|I|=M_{0}}{\mathds{1}}\{\mathcal{E}_{R_{I}}\}={M-|\widehat{I}|\choose M_{0}}$. And, due to the computational constraint, there are at most $\lfloor\frac{B_{\textnormal{train}}}{n}\rfloor$ many rounds $r$ for which $|\mathcal{D}^{(r)}_{\ell}|=n$ and so we must have $|\widehat{I}|\leq\lfloor\frac{B_{\textnormal{train}}}{n}\rfloor$. Combining all our calculations, then,
$\mathbb{P}\{{\widehat{T}}=1\}\leq\frac{\sum_{I\subset[M],|I|=M_{0}}\mathbb{P}\{{\widehat{T}}=1;\mathcal{E}_{R_{I}}\}}{{M-\lfloor\frac{B_{\textnormal{train}}}{n}\rfloor\choose M_{0}}}\leq\frac{{M\choose M_{0}}\cdot\alpha}{{M-\lfloor\frac{B_{\textnormal{train}}}{n}\rfloor\choose M_{0}}}.$
Finally, taking $M\to\infty$ (and setting $M_{0}=1+\lfloor M\cdot\frac{\delta-\delta_{\epsilon}^{*}}{1-\delta_{\epsilon}^{*}}\rfloor\approx M\cdot\frac{\delta-\delta_{\epsilon}^{*}}{1-\delta_{\epsilon}^{*}}$), we have $\lim_{M\to\infty}\frac{{M\choose M_{0}}}{{M-\lfloor\frac{B_{\textnormal{train}}}{n}\rfloor\choose M_{0}}}=\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{\lfloor\frac{B_{\textnormal{train}}}{n}\rfloor}$, which proves that $\mathbb{P}\{{\widehat{T}}=1\}\leq\alpha\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{\frac{B_{\textnormal{train}}}{n}}$.
Proof of Lemma 1.
We define an algorithm $\mathcal{A}^{\prime}$, that is a slight modification of $\mathcal{A}$. The idea will be that $\mathcal{A}^{\prime}$ is unstable, but cannot be distinguished from $\mathcal{A}$ on the event $\mathcal{E}_{R}$. We define [^3] In this construction, we are using the fact that algorithms produce functions that map from $\mathcal{X}$ to $\mathbb{R}$, as in (2)—i.e., fitted models return real-valued predictions. In some settings, for instance if $\mathcal{Y}=\{0,1\}$ or $\mathcal{Y}=[0,1]$, it may be more natural to constrain fitted models to return predictions in a bounded subset of $\mathbb{R}$, e.g., in $[0,1]$. In that type of setting, the same proof techniques for all our results still apply with a slight modification of the construction of the modified algorithm $\mathcal{A}^{\prime}$. The resulting bound on power is slightly weaker, but the difference vanishes as $\epsilon\to 0$.
$\mathcal{A}^{\prime}(\mathcal{D};\xi)=\begin{cases}\mathcal{A}_{1}(\mathcal{D};\xi),&\textnormal{ if }|\mathcal{D}|=n\textnormal{ and }\xi\in R,\\ \mathcal{A}(\mathcal{D};\xi),&\textnormal{ otherwise},\end{cases}$
where $\mathcal{A}_{1}(\mathcal{D};\xi)$ is the function given by $\mathcal{A}_{1}(\mathcal{D};\xi)(x)=1+\epsilon+\max_{\sigma\in S_{|\mathcal{D}|}}\mathcal{A}((\mathcal{D}_{\sigma})_{-|\mathcal{D}|};\xi)(x)$. Here $S_{|\mathcal{D}|}$ denotes the set of all permutations on a set of size $|\mathcal{D}|$, and the notation $(\mathcal{D}_{\sigma})_{-|\mathcal{D}|}$ means that we first permute data set $\mathcal{D}$, i.e., $\left((x_{\sigma(1)},y_{\sigma(1)}),\ldots,(x_{\sigma(|\mathcal{D}|)},y_{\sigma(|\mathcal{D}|)})\right)$, and then remove the last data point. In other words, we have constructed $\mathcal{A}^{\prime}$ so that if the training data set $\mathcal{D}_{n}$ has $n$ data points and $\xi\in R$, then we will inevitably see instability, i.e., we will have $|\mathcal{A}^{\prime}(\mathcal{D}_{n};\xi)(x)-\mathcal{A}^{\prime}(\mathcal{D}_{n-1};\xi)(x)|>\epsilon$, for any $x$. First we verify that $(\mathcal{A}^{\prime},P,n)$ is not stable: we have
$\mathbb{P}_{P}\left\{|\mathcal{A}^{\prime}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}^{\prime}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon\right\}$
$=\mathbb{P}_{P}\left\{|\mathcal{A}^{\prime}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}^{\prime}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon,\xi\in R\right\}$
$\hskip 72.26999pt{}+\mathbb{P}_{P}\left\{|\mathcal{A}^{\prime}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}^{\prime}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon,\xi\notin R\right\}$
$=\mathbb{P}_{P}\left\{\xi\in R\right\}+\mathbb{P}_{P}\left\{|\mathcal{A}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon,\xi\notin R\right\},$
since by construction of $\mathcal{A}^{\prime}$, if $\xi\in R$ then we must have instability; while if $\xi\not\in R$ then $\mathcal{A}^{\prime}$ returns the same fitted model as $\mathcal{A}$. Moreover, since $\xi\sim\textnormal{Unif}[0,1]$, we have $\mathbb{P}_{P}\left\{\xi\in R\right\}=\textnormal{Leb}(R)$. And, by definition of the function $f$, $\mathbb{P}_{P}\left\{|\mathcal{A}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon\mid\xi\not\in R\right\}=\mathbb{E}_{P}[f(\xi)\mid\xi\not\in R]$. Combining everything, then,
$\mathbb{P}_{P}\left\{|\mathcal{A}^{\prime}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}^{\prime}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon\right\}=\textnormal{Leb}(R)+(1-\textnormal{Leb}(R))\cdot\mathbb{E}_{P}[f(\xi)\mid\xi\not\in R].$
Since this quantity is assumed to be $>\delta$, by the assumption in the lemma, this verifies that $(\mathcal{A}^{\prime},P,n)$ is not $(\epsilon,\delta)$-stable. Thus by validity of ${\widehat{T}}$ (3) we have $\mathbb{P}_{P}({\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{\ell},\mathcal{D}_{u})=1)\leq\alpha$. Next, observe that when we run the test ${\widehat{T}}$ on data sets $\mathcal{D}_{\ell}$ and $\mathcal{D}_{u}$ with either algorithm $\mathcal{A}$ or $\mathcal{A}^{\prime}$, by construction of $\mathcal{A}^{\prime}$ the outcome of the test must be the same on the event $\mathcal{E}_{R}$. So, we have
$\mathbb{P}_{P}({\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})=1;\mathcal{E}_{R})=\mathbb{P}_{P}({\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{\ell},\mathcal{D}_{u})=1;\mathcal{E}_{R})\leq\mathbb{P}_{P}({\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{\ell},\mathcal{D}_{u})=1)\leq\alpha.$
5.2 Statistical limit: bounding power due to limited $Y$ data
For any $y\in\mathcal{Y}$, define $\mathcal{E}_{y}$ as the event that $y\not\in\mathcal{D}^{(1)}_{\ell}\cup\dots\cup\mathcal{D}^{({\widehat{r}})}_{\ell}$, when running the test ${\widehat{T}}$. The key step for this part of the proof is the following lemma:
Lemma 2.
In the setting of Theorem 1, for any $y\in\mathcal{Y}$, $\mathbb{P}\{{\widehat{T}}=1;\mathcal{E}_{y}\}\leq\alpha\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{\frac{N_{\ell}}{n}}$.
The bound on power is then implied by this lemma, as follows. First, we can assume $|\mathcal{Y}|>B_{\textnormal{train}}$, since otherwise the bound is trivial. Fix any integer $M$ with $B_{\textnormal{train}}<M\leq|\mathcal{Y}|$ (note that $|\mathcal{Y}|$ may be finite or infinite), and let $y_{1},\dots,y_{M}\in\mathcal{Y}$ be distinct. We then have
$\sum_{i=1}^{M}\mathbb{P}\{{\widehat{T}}=1;\mathcal{E}_{y_{i}}\}=\mathbb{E}\left[\sum_{i=1}^{M}{\mathds{1}}\{{\widehat{T}}=1;\mathcal{E}_{y_{i}}\}\right]=\mathbb{E}\left[{\mathds{1}}\{{\widehat{T}}=1\}\cdot\sum_{i=1}^{M}{\mathds{1}}\{\mathcal{E}_{y_{i}}\}\right]\\ \geq\mathbb{E}\left[{\mathds{1}}\{{\widehat{T}}=1\}\cdot(M-B_{\textnormal{train}})\right]=\mathbb{P}\{{\widehat{T}}=1\}\cdot(M-B_{\textnormal{train}}),$
where the inequality holds since, due to the computational constraint, $y\in\mathcal{D}^{(1)}_{\ell}\cup\dots\cup\mathcal{D}^{({\widehat{r}})}_{\ell}$ can hold for at most $B_{\textnormal{train}}$ many values $y\in\mathcal{Y}$—and therefore, at least $M-B_{\textnormal{train}}$ many of the events $\mathcal{E}_{y_{1}},\dots,\mathcal{E}_{y_{M}}$ must hold. Therefore,
$\mathbb{P}\{{\widehat{T}}=1\}\leq\frac{\sum_{i=1}^{M}\mathbb{P}\{{\widehat{T}}=1;\mathcal{E}_{y_{i}}\}}{M-B_{\textnormal{train}}}\leq\frac{M\cdot\alpha\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{N_{\ell}/n}}{M-B_{\textnormal{train}}},$
where the last step applies Lemma 2. Since this holds for any $M$ with $B_{\textnormal{train}}<M\leq|\mathcal{Y}|$, we have
$\mathbb{P}\{{\widehat{T}}=1\}\leq\alpha\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{N_{\ell}/n}\cdot\inf_{M\,:\,B_{\textnormal{train}}<M\leq|\mathcal{Y}|}\frac{M}{M-B_{\textnormal{train}}}.$
If $|\mathcal{Y}|<\infty$, then $\inf_{M\,:\,B_{\textnormal{train}}<M\leq|\mathcal{Y}|}\frac{M}{M-B_{\textnormal{train}}}=\frac{|\mathcal{Y}|}{|\mathcal{Y}|-B_{\textnormal{train}}}$. If instead $|\mathcal{Y}|=\infty$, then $\inf_{M\,:\,B_{\textnormal{train}}<M\leq|\mathcal{Y}|}\frac{M}{M-B_{\textnormal{train}}}=1$. For both cases, then, we have obtained the desired result.
Proof of Lemma 2.
At a high level, the main idea of the proof of this part is to construct a corrupted distribution $P^{\prime}$, and a modified algorithm $\mathcal{A}^{\prime}$, so that $(\mathcal{A}^{\prime},P^{\prime},n)$ is not stable, but it is difficult to distinguish between $\mathcal{A}^{\prime},P^{\prime}$ and $\mathcal{A},P$ given limited data. This proof extends the ideas constructed in the proof of (Kim and Barber,, 2023, Theorem 2) to this more complex setting where $|\mathcal{Y}|$ is no longer assumed to be infinite.
Fix any $c\in[0,1]$. Define a new distribution
$P^{\prime}=c\cdot(P_{X}\times\delta_{y})+(1-c)\cdot P,$
where $P_{X}$ is the marginal of $X$ under $(X,Y)\sim P$, while $\delta_{y}$ is the point mass at $y$—note that the marginal distribution of $X$ is the same under $P^{\prime}$ as under $P$. Following the construction in the proof of (Kim and Barber,, 2023, Theorem 2), we define also a new algorithm $\mathcal{A}^{\prime}$ as
$\mathcal{A}^{\prime}(\mathcal{D};\xi)=\begin{cases}\mathcal{A}_{1}(\mathcal{D};\xi),&\textnormal{ if }|\mathcal{D}|=n\textnormal{ and }y\in\mathcal{D},\\ \mathcal{A}(\mathcal{D};\xi),&\textnormal{ otherwise},\end{cases}$
where the informal notation $y\in\mathcal{D}$ denotes that at least one of the response values in the data points contained in $\mathcal{D}$ is equal to $y$, and where $\mathcal{A}_{1}$ is defined as in the proof of Lemma 1.
Next we verify that $(\mathcal{A}^{\prime},P^{\prime},n)$ is not stable. Let $\mathcal{D}_{n}$ denote a training set of size $n$ drawn i.i.d. from $P^{\prime}$. Let $\mathcal{E}_{\mathcal{D}_{n}\sim P}$ be the event that, when sampling the training set $\mathcal{D}_{n}$, all $n$ samples are drawn from the second component of the mixture, i.e., from $P$. Then this event holds with probability $(1-c)^{n}$ by construction. Moreover, on the event that $\mathcal{E}_{\mathcal{D}_{n}\sim P}$ does not occur, we have
$\mathbb{P}_{P^{\prime}}\{|\mathcal{A}^{\prime}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}^{\prime}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon\mid\mathcal{E}_{\mathcal{D}_{n}\sim P}^{c}\}=1$
(since, if at least one training sample is drawn from the first component of the mixture, i.e., $P_{X}\times\delta_{y}$, then we have $y\in\mathcal{D}_{n}$ and so instability must occur by construction of $\mathcal{A}^{\prime}$), while if instead $\mathcal{E}_{\mathcal{D}_{n}\sim P}$ does occur, we calculate
$\mathbb{P}_{P^{\prime}}\{|\mathcal{A}^{\prime}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}^{\prime}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon\mid\mathcal{E}_{\mathcal{D}_{n}\sim P}\}\\ \geq\mathbb{P}_{P}\{|\mathcal{A}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon\}=\delta^{*}_{\epsilon},$
where the inequality holds by construction of $\mathcal{A}^{\prime}$. Therefore,
$\mathbb{P}_{P^{\prime}}\{|\mathcal{A}^{\prime}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}^{\prime}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon\}\geq\left[1-(1-c)^{n}\right]+(1-c)^{n}\delta^{*}_{\epsilon},$
and in particular, this quantity is $>\delta$ for any $c>1-\left(\frac{1-\delta}{1-\delta^{*}_{\epsilon}}\right)^{1/n}$. Finally, since $\mathcal{A}^{\prime}$ returns identical output to $\mathcal{A}$ on each training set $\mathcal{D}^{(r)}_{\ell}$ as long as $y\not\in\mathcal{D}^{(r)}_{\ell}$, therefore on the event $\mathcal{E}_{y}$ (i.e., the event $y\not\in\mathcal{D}^{(1)}_{\ell}\cup\dots\cup\mathcal{D}^{({\widehat{r}})}_{\ell}$) we must have ${\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{\ell},\mathcal{D}_{u})=1$ if and only if ${\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})=1$ (a more formal proof of this type of statement can be found in the proof of Kim and Barber, (2023, Theorem 2)). Combining these calculations, we have
$\mathbb{P}_{P}\{{\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})=1;\mathcal{E}_{y}\}=\mathbb{P}_{P}\{{\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{\ell},\mathcal{D}_{u})=1;\mathcal{E}_{y}\}\leq\mathbb{P}_{P}\{{\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{\ell},\mathcal{D}_{u})=1\}\\ \leq(1-c)^{-N_{\ell}}\cdot\mathbb{P}_{P^{\prime}}\{{\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{\ell},\mathcal{D}_{u})=1\}\leq\alpha(1-c)^{-N_{\ell}},$
where the last step holds by validity of ${\widehat{T}}$ (3), since $(\mathcal{A}^{\prime},P^{\prime},n)$ is not $(\epsilon,\delta)$-stable. We have established that this is true for any $c>1-\left(\frac{1-\delta}{1-\delta^{*}_{\epsilon}}\right)^{1/n}$, and therefore,
$\mathbb{P}_{P}\{{\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})=1;\mathcal{E}_{y}\}\leq\inf_{c>1-\left(\frac{1-\delta}{1-\delta^{*}_{\epsilon}}\right)^{1/n}}\alpha(1-c)^{-N_{\ell}}=\alpha\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{N_{\ell}/n}.$
5.3 Statistical limit: bounding power due to limited $X$ data
Analogously to our proof for limited $Y$ data, for any $x\in\mathcal{X}$ we define $\mathcal{E}_{x}$ as the event that $x\not\in\mathcal{D}^{(1)}_{\ell}\cup\dots\cup\mathcal{D}^{({\widehat{r}})}_{\ell}$, when running the test ${\widehat{T}}$. For the third bound on power, the key step is to show the following result:
Lemma 3.
In the setting of Theorem 1, if $\delta(1+\frac{1}{en})<1$, then for any $x\in\mathcal{X}$, $\mathbb{P}\{{\widehat{T}}=1;\mathcal{E}_{x}\}\leq\alpha\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta(1+\frac{1}{en})}\right)^{\frac{N_{\ell}+N_{u}}{n}}.$
With this lemma in place, the proof for the power upper bound due to limited data in $\mathcal{X}$ space is exactly the same as for the previous step (i.e., with $Y$ in place of $X$), so we omit the details and simply turn to the proof of the lemma.
Proof of Lemma 3.
The proof of this lemma follows a similar construction as for the proof of Lemma 2 (for $Y$ in place of $X$), but is more complex since the value of $X$ plays a role both in the training data and in the test point.
Fix any $c\in[0,1]$. Define a new distribution
$P^{\prime}=c\cdot(\delta_{x}\times P_{Y})+(1-c)\cdot P,$
where $P_{Y}$ is the marginal of $Y$ under $(X,Y)\sim P$, while $\delta_{x}$ is the point mass at $x$. Now define a new algorithm $\mathcal{A}^{\prime}$ as
$\mathcal{A}^{\prime}(\mathcal{D};\xi)=\begin{cases}\mathcal{A}_{1}(\mathcal{D};\xi),&\textnormal{ if }|\mathcal{D}|=n\textnormal{ and }x\in\mathcal{D},\\ \mathcal{A}(\mathcal{D};\xi),&\textnormal{ otherwise},\end{cases}$ (6)
where $\mathcal{A}_{1}$ is defined as in the proof of Lemma 1, and where $x\in\mathcal{D}$ is interpreted as before. Now we see that, for an appropriate choice of $c$, $(\mathcal{A}^{\prime},P^{\prime},n)$ is unstable. We have
$\mathbb{P}_{P^{\prime}}\left\{|\mathcal{A}^{\prime}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}^{\prime}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon\right\}\geq\left[1-(1-c)^{n}\right]+(1-c)^{n+1}\delta^{*}_{\epsilon},$
where, similarly to the analogous calculation in the proof of Lemma 2, the term in square brackets arises from the event that at least one of the $n$ training points is drawn from the first component in the mixture distribution $P^{\prime}$, i.e., from $\delta_{x}\times P_{Y}$, while the remaining term arises from the event that all $n$ training points and the test point $X_{n+1}$ are drawn from the second component, $P$. Next, we have
$1-(1-c)^{n}+(1-c)^{n+1}\delta^{*}_{\epsilon}\geq 1-(1-c)^{n}(1-\delta^{*}_{\epsilon})-\frac{\delta^{*}_{\epsilon}}{en}\geq 1-(1-c)^{n}(1-\delta^{*}_{\epsilon})-\frac{\delta}{en},$
where the first inequality holds since $((1-c)^{n}-(1-c)^{n+1})\leq\sup_{t\in[0,1]}\{t^{n}-t^{n+1}\}\leq\frac{1}{en}$, and the second inequality holds since $\delta^{*}_{\epsilon}\leq\delta$. Taking any $c>1-\left(\frac{1-\delta(1+\frac{1}{en})}{1-\delta^{*}_{\epsilon}}\right)^{1/n}$, we have $\mathbb{P}_{P^{\prime}}\left\{|\mathcal{A}^{\prime}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}^{\prime}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon\right\}>\delta$.
Finally, since $\mathcal{A}^{\prime}$ returns identical output to $\mathcal{A}$ on each training set $\mathcal{D}^{(r)}_{\ell}$ as long as $x\not\in\mathcal{D}^{(r)}_{\ell}$, on the event $\mathcal{E}_{x}$ we must have ${\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{\ell},\mathcal{D}_{u})=1$ if and only if ${\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})=1$. Then
$\mathbb{P}_{P}\{{\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})=1;\mathcal{E}_{x}\}=\mathbb{P}_{P}\{{\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{\ell},\mathcal{D}_{u})=1;\mathcal{E}_{x}\}\leq\mathbb{P}_{P}\{{\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{\ell},\mathcal{D}_{u})=1\}\\ \leq(1-c)^{-(N_{\ell}+N_{u})}\cdot\mathbb{P}_{P^{\prime}}\{{\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{\ell},\mathcal{D}_{u})=1\}\leq\alpha(1-c)^{-(N_{\ell}+N_{u})},$
where the last step holds by validity of ${\widehat{T}}$ (3), since $(\mathcal{A}^{\prime},P^{\prime},n)$ is not $(\epsilon,\delta)$-stable. We have established that this is true for any $c\in[0,1]$ with $c>1-\left(\frac{1-\delta(1+\frac{1}{en})}{1-\delta^{*}_{\epsilon}}\right)^{1/n}$, this proves the desired bound. ∎
Acknowledgements
R.F.B. was supported by the Office of Naval Research via grant N00014-20-1-2337 and by the National Science Foundation via grant DMS-2023109.
Appendix A Formal definition of a black-box test with computational constraints
In this section, we give a formal definition of a black-box test under a computational constraint defined by budget $B_{\textnormal{train}}$. This definition is visualized in Figure 1.
Let $\mathfrak{D}_{\ell}=\cup_{m\geq 0}(\mathcal{X}\times\mathcal{Y})^{m}$ denotes the space of labeled data sets of any size, and let $\mathfrak{D}_{u}=\cup_{m\geq 0}(\mathcal{X})^{m}$ denotes the space of unlabeled data sets of any size.
Definition 2 (Black-box test under computational constraints).
Consider any function ${\widehat{T}}$ that takes as input an algorithm $\mathcal{A}$, a labeled data set $\mathcal{D}_{\ell}\in\mathfrak{D}_{\ell}$, and an unlabeled data set $\mathcal{D}_{u}\in\mathfrak{D}_{u}$, and returns a (possibly randomized) output ${\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})\in\{0,1\}$. We say that ${\widehat{T}}$ is a black-box test with computational budget $B_{\textnormal{train}}>0$ for model training if it can be defined as follows: for some functions $g^{(1)},g^{(2)},\dots$ and $g$, and for random seeds, $\zeta^{(1)},\zeta^{(2)},\dots,\zeta\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}\textnormal{Unif}[0,1]$,
  1. 1.
    For each stage $r=1,2,3,\ldots$,
    1. (a)
      If $r=1$, generate a new labeled data set and a new random seed,
      $(\mathcal{D}_{\ell}^{(1)},\xi^{(1)})=g^{(1)}(\mathcal{D}_{\ell},\mathcal{D}_{u},\zeta^{(1)}).$
      • If some stopping criterion is reached or $|\mathcal{D}^{(1)}_{\ell}|>B_{\textnormal{train}}$, exit the for loop; otherwise compute the fitted model ${\widehat{f}}^{(1)}=\mathcal{A}(\mathcal{D}_{\ell}^{(1)};\xi^{(1)})$.
    2. (b)
      If $r\geq 2$, generate a labeled data set and a new random seed,
      $(\mathcal{D}_{\ell}^{(r)},\xi^{(r)})=g^{(r)}\big{(}\mathcal{D}_{\ell},\mathcal{D}_{u},(\mathcal{D}_{\ell}^{(s)})_{1\leq s<r},({\widehat{f}}^{(s)})_{1\leq s<r},(\zeta^{(s)})_{1\leq s\leq r},(\xi^{(s)})_{1\leq s<r}\big{)}.$
      • If some stopping criterion is reached or $\sum_{j=1}^{r}|\mathcal{D}_{\ell}^{(j)}|>B_{\textnormal{train}}$, exit the for loop; otherwise compute the fitted model ${\widehat{f}}^{(r)}=\mathcal{A}(\mathcal{D}_{\ell}^{(r)};\xi^{(r)})$.
  2. 2.
    Let ${\widehat{r}}$ denote the total number of rounds when stopping.
  3. 3.
    Finally, return
    ${\widehat{T}}=g\big{(}\mathcal{D}_{\ell},\mathcal{D}_{u},(\mathcal{D}_{\ell}^{(r)})_{1\leq r\leq{\widehat{r}}},({\widehat{f}}^{(r)})_{1\leq r\leq{\widehat{r}}},(\zeta^{(r)})_{1\leq r\leq{\widehat{r}}},(\xi^{(r)})_{1\leq r\leq{\widehat{r}}},\zeta\big{)}.$
The random seeds $\zeta^{(1)},\zeta^{(2)},\dots$ are included to allow for randomization at each stage of the test, if desired. For example, if the data set $\mathcal{D}^{(r)}_{\ell}$ is computed via subsampling from the available real data $\mathcal{D}_{\ell}$ (e.g., a bootstrap sample), then $\zeta^{(r)}$ may be used to generate the random subsample. Similarly, the random seed $\zeta$ is included to allow for randomization, if desired, in the final step of the test—i.e., computing ${\widehat{T}}\in\{0,1\}$.
Appendix B Black-box tests with black-box models
We next turn to a different setting, where both the algorithm $\mathcal{A}$ and its outputs (the fitted models ${\widehat{f}}$) are treated as black-box. To motivate this setting, consider an algorithm $\mathcal{A}$ returning a fitted model ${\widehat{f}}$ that is given by neural network with many layers. Our ability to study the behavior of this complex ${\widehat{f}}$ analytically is likely to be limited—for example, it is known that computing the Lipschitz constant of a deep neural network is hard (Virmaux and Scaman,, 2018). In many settings we may only be able to study properties of the fitted model ${\widehat{f}}$ empirically, i.e., by sampling a large number of values of $X$ and evaluating ${\widehat{f}}$ on these values. (In contrast, if $\mathcal{A}$ is a complex variable selection procedure but returns only simple models ${\widehat{f}}$, such as linear models, then we would be able to study ${\widehat{f}}$ analytically—the transparent-model setting that we have considered up until now, in this paper.)
To allow for the setting of black-box models, we now modify our definition of a black-box test, as follows. To help compare to Definition 2, we use boldface text to highlight the parts of the definition that have been modified.
Definition 3 (Black-box test with black-box models, under computational constraints).
Consider any function ${\widehat{T}}$ that takes as input an algorithm $\mathcal{A}$, a labeled data set $\mathcal{D}_{\ell}\in\mathfrak{D}_{\ell}$, and an unlabeled data set $\mathcal{D}_{u}\in\mathfrak{D}_{u}$, and returns a (possibly randomized) output ${\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})\in\{0,1\}$. We say that ${\widehat{T}}$ is a black-box test, with computational budgets $B_{\textnormal{train}}>0$ for model training and $\boldsymbol{B_{\textnormal{eval}}>0}$ for model evaluation, if it can be defined as follows: for some functions $g^{(1)},g^{(2)},\dots$ and $g$, and for random seeds, $\zeta^{(1)},\zeta^{(2)},\dots,\zeta\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}\textnormal{Unif}[0,1]$,
  1. 1.
    For each stage $r=1,2,3,\ldots$,
    1. (a)
      If $r=1$, generate a new labeled data set, an unlabeled data set, and a new random seed,
      $(\mathcal{D}_{\ell}^{(1)},\boldsymbol{\mathcal{X}^{(1)}},\xi^{(1)})=g^{(1)}(\mathcal{D}_{\ell},\mathcal{D}_{u},\zeta^{(1)}).$
      • If some stopping criterion is reached or $|\mathcal{D}^{(1)}_{\ell}|>B_{\textnormal{train}}$ or $\boldsymbol{|\mathcal{X}^{(1)}|>B_{\textnormal{eval}}}$, exit the for loop; otherwise compute the fitted model ${\widehat{f}}^{(1)}=\mathcal{A}(\mathcal{D}_{\ell}^{(1)};\xi^{(1)})$, and evaluate the fitted model by computing $\boldsymbol{{\widehat{f}}^{(1)}(\mathcal{X}^{(1)})}$ [^4] Here ${\widehat{f}}^{(1)}(\mathcal{X}^{(1)})$ is understood in the pointwise sense, i.e., we compute ${\widehat{f}}^{(1)}(x)$ for each $x\in\mathcal{X}^{(1)}$.
    2. (b)
      If $r\geq 2$, generate a labeled data set, an unlabeled data set, and a new random seed,
      $(\mathcal{D}_{\ell}^{(r)},\boldsymbol{\mathcal{X}^{(r)}},\xi^{(r)})\\ =g^{(r)}\big{(}\mathcal{D}_{\ell},\mathcal{D}_{u},(\mathcal{D}_{\ell}^{(s)})_{1\leq s<r},\boldsymbol{(\mathcal{X}^{(s)})_{1\leq s<r},({\widehat{f}}^{(s)}(\mathcal{X}^{(s)}))_{1\leq s<r}},\\ (\zeta^{(s)})_{1\leq s\leq r},(\xi^{(s)})_{1\leq s<r}\big{)}.$
      • If some stopping criterion is reached or $\sum_{j=1}^{r}|\mathcal{D}_{\ell}^{(j)}|>B_{\textnormal{train}}$ or $\boldsymbol{\sum_{j=1}^{r}|\mathcal{X}^{(j)}|>B_{\textnormal{eval}}}$, exit the for loop; otherwise compute the fitted model ${\widehat{f}}^{(r)}=\mathcal{A}(\mathcal{D}_{\ell}^{(r)};\xi^{(r)})$, and evaluate the fitted model by computing $\boldsymbol{{\widehat{f}}^{(r)}(\mathcal{X}^{(r)})}$.
  2. 2.
    Let ${\widehat{r}}$ denote the total number of rounds when stopping.
  3. 3.
    Finally, return
    ${\widehat{T}}=g\big{(}\mathcal{D}_{\ell},\mathcal{D}_{u},(\mathcal{D}_{\ell}^{(r)})_{1\leq r\leq{\widehat{r}}},\boldsymbol{(\mathcal{X}^{(r)})_{1\leq r\leq{\widehat{r}}}},\boldsymbol{({\widehat{f}}^{(r)}(\mathcal{X}^{(r)}))_{1\leq r\leq{\widehat{r}}}},(\zeta^{(r)})_{1\leq r\leq{\widehat{r}}},(\xi^{(r)})_{1\leq r\leq{\widehat{r}}},\zeta\big{)}.$
A pictorial illustration of a black-box test with black-box models is given in Figure 2.
The main result for the power upper bound of testing algorithmic stability via black-box test with black-box models is given as follows.
Theorem 2.
Fix any $\epsilon\geq 0$, $\delta\in[0,1)$. Let ${\widehat{T}}$ be any black-box test with black-box models, as in Definition 3. Suppose ${\widehat{T}}$ has computational budgets $B_{\textnormal{train}}$ for model training and $B_{\textnormal{eval}}$ for model evaluation, and satisfies the assumption-free validity condition (3) at level $\alpha$. Then for any $(\mathcal{A},P,n)$ that is $(\epsilon,\delta)$-stable (i.e. $\delta^{*}_{\epsilon}\leq\delta$ and the alternative holds), the power of ${\widehat{T}}$ is bounded as
[…]
Comparison to existing work.
Similarly to the transparent-model setting, we can compare to the results of Kim and Barber, (2023) in the case of infinite data spaces $\mathcal{X}$ and $\mathcal{Y}$. When $|\mathcal{X}|=|\mathcal{Y}|=\infty$, then even for arbitrarily large or unlimited $B_{\textnormal{train}}$ and $B_{\textnormal{eval}}$ (i.e., we may make arbitrarily many—but still finitely many—calls to train $\mathcal{A}$ and to evaluate the fitted models), we have $\frac{B_{\textnormal{train}}}{|\mathcal{Y}|},\frac{B_{\textnormal{train}}+B_{\textnormal{eval}}}{|\mathcal{X}|}\to 0$ and so (7) implies
$\mathbb{P}\{{\widehat{T}}=1\}\leq\alpha\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{\frac{N_{\ell}}{n}\wedge\frac{N_{\ell}+N_{u}}{n+1}}.$
This exactly recovers the result of Kim and Barber, (2023, Theorem 2) and demonstrates the hardness of testing algorithmic stability when the available data is limited. At the same time, it has been shown in Theorem 1 of Kim and Barber, (2023) that this upper bound on power can be achieved (under some mild assumptions) by a simple Binomial test based on sample splitting.
Input: labeled data $\mathcal{D}_{\ell}$, unlabeled data $\mathcal{D}_{u}$, algorithm $\mathcal{A}$, seeds $\zeta,\zeta_{1},\zeta_{2},\dots$ Iterate for rounds $r=1,2,\ldots,{\widehat{r}}$, subject to computational constraints $\sum_{r=1}^{{\widehat{r}}}|\mathcal{D}^{(r)}_{\ell}|\leq B_{\textnormal{train}}$ and $\sum_{j=1}^{{\widehat{r}}}|\mathcal{X}^{(r)}|\leq B_{\textnormal{eval}}$ Generate labeled data set $\mathcal{D}^{(r)}_{\ell}$, random seed $\xi^{(r)}$ and an unlabeled data set, $\mathcal{X}^{(r)}$ Call the algorithm to train a model: ${\widehat{f}}^{(r)}=\mathcal{A}(\mathcal{D}^{(r)}_{\ell};\xi^{(r)})$, then evaluate ${\widehat{f}}^{(r)}(\mathcal{X}^{(r)})$ Output: ${\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})\in\{0,1\}$
Figure 2: Black-box test with black-box models under computational constraints. Here in round $r$, the generating process of the new labeled data set $\mathcal{D}_{\ell}^{(r)}$ for training, the unlabeled data set $\mathcal{X}^{(r)}$ for evaluation and the new random seed $\xi^{(r)}$ can depend on the input data and on all the past information from previous rounds—e.g., we may create a new labeled data set and an unlabeled data set by resampling from past data. The input random seeds $\zeta,\zeta_{1},\zeta_{2},\dots\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}\textnormal{Unif}[0,1]$ may be used to provide randomization if desired. The procedure stops at a data-dependent stopping time ${\widehat{r}}$ (i.e., we may use any stopping criterion), subject to the computational constraints that (1) the total number of training points used in calls to $\mathcal{A}$ cannot exceed the budget $B_{\textnormal{train}}$ and (2) the total number of data points used to evaluate the fitted models cannot exceed the budget $B_{\textnormal{eval}}$.
B.1 Proof of Theorem 2
The first three terms in the power bound are the same as the ones appearing in Theorem 1. Since this new result, Theorem 2, places strictly more constraints on the test ${\widehat{T}}$ (i.e., due to the model evaluation budget $B_{\textnormal{train}}$), these three bounds on power hold in this setting as well. To complete the proof, then, we only need to establish the last bound, namely,
$\mathbb{P}\{{\widehat{T}}=1\}\leq\alpha\cdot\frac{\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{\frac{N_{\ell}+N_{u}}{n+1}}}{1-\frac{B_{\textnormal{train}}+B_{\textnormal{eval}}}{|\mathcal{X}|}\wedge 1}.$
The proof is similar to the one in Section 5.3, except now we define the event ~$\widetilde{\mathcal{E}}_{x}$ as $\mathcal{E}_{x}=\mathcal{D}_{\ell}^{(1)}\cup\mathcal{X}^{(1)}\cup\dots\cup\mathcal{D}_{\ell}^{({\widehat{r}})}\cup\mathcal{X}^{({\widehat{r}})}$ since we are now interested in the role of limited $X$ data in terms of model evaluation as well as model training. As before, we will need a lemma:
Lemma 4.
In the setting of Theorem 2, for any $x\in\mathcal{X}$,
$\mathbb{P}\{{\widehat{T}}=1;\widetilde{\mathcal{E}}_{x}\}\leq\alpha\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{\frac{N_{\ell}+N_{u}}{n+1}}.$
Now we prove that this lemma implies the desired bound on power. This calculation is similar as for the transparent-model setting (as in Section 5.2).
First, we can assume $|\mathcal{X}|>B_{\textnormal{train}}+B_{\textnormal{eval}}$, since otherwise the bound is trivial. Fix any integer $M$ with $B_{\textnormal{train}}+B_{\textnormal{eval}}<M\leq|\mathcal{X}|$ (note that $|\mathcal{X}|$ may be finite or infinite), and let $x_{1},\dots,x_{M}\in\mathcal{X}$ be distinct. We then have
$\sum_{i=1}^{M}\mathbb{P}\{{\widehat{T}}=1;\widetilde{\mathcal{E}}_{x_{i}}\}=\mathbb{E}\left[{\mathds{1}}\{{\widehat{T}}=1\}\cdot\sum_{i=1}^{M}{\mathds{1}}\{\widetilde{\mathcal{E}}_{x_{i}}\}\right]\\ \geq\mathbb{E}\left[{\mathds{1}}\{{\widehat{T}}=1\}\cdot(M-B_{\textnormal{train}}-B_{\textnormal{eval}})\right]=\mathbb{P}\{{\widehat{T}}=1\}\cdot(M-B_{\textnormal{train}}-B_{\textnormal{eval}}),$
where the inequality holds since, due to the computational constraint, $x\in\mathcal{D}^{(1)}_{\ell}\cup\mathcal{X}^{(1)}\cup\dots\cup\mathcal{D}^{({\widehat{r}})}_{\ell}\cup\mathcal{X}^{({\widehat{r}})}$ can hold for at most $B_{\textnormal{train}}+B_{\textnormal{eval}}$ many values $x\in\mathcal{X}$. Therefore,
$\mathbb{P}\{{\widehat{T}}=1\}\leq\frac{\sum_{i=1}^{M}\mathbb{P}\{{\widehat{T}}=1;\widetilde{\mathcal{E}}_{x_{i}}\}}{M-B_{\textnormal{train}}-B_{\textnormal{eval}}}\leq\frac{M\cdot\alpha\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{\frac{N_{\ell}+N_{u}}{n+1}}}{M-B_{\textnormal{train}}-B_{\textnormal{eval}}},$
where the last step holds by Lemma 4. Since this holds for any $M$ with $B_{\textnormal{train}}-B_{\textnormal{eval}}<M\leq|\mathcal{X}|$, as in Section 5.2 we can take $M=|\mathcal{X}|$ to complete the proof for the case $|\mathcal{X}|<\infty$, or $M\to\infty$ for the case $|\mathcal{X}|=\infty$.
Proof of Lemma 4.
The proof of this lemma is similar to that of Lemma 3 but now we must also account for limits on model evaluation. Fix any $c\in[0,1]$. We define $P^{\prime}$ exactly as in the proof of Lemma 3, i.e., $P^{\prime}=c\cdot(\delta_{x}\times P_{Y})+(1-c)\cdot P$, but the modified algorithm is defined differently. We define $\mathcal{A}^{\prime}$ by specifying, for any training set $\mathcal{D}$ and random seed $\xi$, the value of the fitted model when evaluated at a test point $x^{\prime}$:
$\mathcal{A}^{\prime}(\mathcal{D};\xi)(x^{\prime})=\begin{cases}\mathcal{A}_{1}(\mathcal{D};\xi)(x^{\prime})&\textnormal{ if }|\mathcal{D}|=n,\textnormal{ and either $x\in\mathcal{D}$ or $x^{\prime}=x$},\\ \mathcal{A}(\mathcal{D};\xi)(x^{\prime}),&\textnormal{ otherwise},\end{cases}$ (8)
where $\mathcal{A}_{1}$ is defined as in the proof of Lemma 1. With this construction, we then have
$\mathbb{P}_{P^{\prime}}\left\{|\mathcal{A}^{\prime}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}^{\prime}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon\right\}\geq\left[1-(1-c)^{n+1}\right]+(1-c)^{n+1}\delta^{*}_{\epsilon},$
where, similarly to the analogous calculation in the proofs of Lemmas 2 and 3, the term in square brackets arises from the event that at least one of the $n+1$ data points (the training set $\mathcal{D}_{n}$ and the test point $X_{n+1}$) is drawn from the first component in the mixture distribution $P^{\prime}$, i.e., from $\delta_{x}\times P_{Y}$, while the remaining term arises from the event that all $n+1$ data points are drawn from the second component, $P$. In particular, we have
$\mathbb{P}_{P^{\prime}}\left\{|\mathcal{A}^{\prime}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}^{\prime}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon\right\}>\delta$
as long as $c>1-\left(\frac{1-\delta}{1-\delta^{*}_{\epsilon}}\right)^{1/(n+1)}$.
Next, since $\mathcal{A}^{\prime}(\mathcal{D}^{(r)}_{\ell};\xi^{(r)})(x^{\prime})=\mathcal{A}^{\prime}(\mathcal{D}^{(r)}_{\ell};\xi^{(r)})(x^{\prime})$ for all $x^{\prime}\in\mathcal{X}^{(r)}$, as long as $x\not\in\mathcal{D}^{(r)}_{\ell}$ and $x^{\prime}\neq x$, this means that running the test ${\widehat{T}}$ with $\mathcal{A}^{\prime}$ or with $\mathcal{A}$ returns identical answers on the event ~$\widetilde{\mathcal{E}}_{x}$. We therefore have
$\mathbb{P}_{P}\{{\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})=1;\widetilde{\mathcal{E}}_{x}\}=\mathbb{P}_{P}\{{\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{\ell},\mathcal{D}_{u})=1;\widetilde{\mathcal{E}}_{x}\}\leq\mathbb{P}_{P}\{{\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{\ell},\mathcal{D}_{u})=1\}\\ \leq(1-c)^{-(N_{\ell}+N_{u})}\cdot\mathbb{P}_{P^{\prime}}\{{\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{\ell},\mathcal{D}_{u})=1\}\leq\alpha(1-c)^{-(N_{\ell}+N_{u})},$
where the last step holds by validity of ${\widehat{T}}$ (3), since $(\mathcal{A}^{\prime},P^{\prime},n)$ is not $(\epsilon,\delta)$-stable. We have established that this is true for any $c\in[0,1]$ with $c>1-\left(\frac{1-\delta}{1-\delta^{*}_{\epsilon}}\right)^{1/(n+1)}$, and so we have
$\mathbb{P}_{P}\{{\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})=1;\widetilde{\mathcal{E}}_{x}\}\leq\alpha\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{\frac{N_{\ell}+N_{u}}{n+1}}.$